# GAD-Generative Learning for HD Map-Free Autonomous Driving

### v2

@@ 3.4 Prediction
We follow a multi-stage design for scene and agent prediction. In the first stage, the pipeline aims to predict the scene level grid map, also known as the heat map or occupancy. We regress a grid map $G_{ij}^{t}=dec(f_{ras})$ with a resolution matches that of the rasterization input by decoding the middle layer features.
Figure 3: Multi-stage Prediction
The decoder, as mentioned in Sec. 3.2, shared an identical design as that of for cost volume but with the supervision from the rasterization of the agents’ future ground truth box. This grid is trained with pixel-wise focal loss inspired from [15]:
$\begin{array}[]{r}L_{occ\_pred}=-\sum_{ij}[\tilde{g_{ij}}(1-\hat{g_{ij}})^{2}log(\hat{g_{ij}})+\\ (1-\tilde{g_{ij}})\hat{g_{ij}}^{2}log(1-\hat{g_{ij}})]\end{array}$ (1)
Here, $\tilde{g_{ij}}$ represents the class label for the grid, where $\tilde{g_{ij}}=1$ indicates that it is occupied.
In the second stage, the model forecasts agent-level trajectories with the scene prior from the predicted grid. For the agent-level prediction, we follow the interactive design. Since the output trajectories are mainly adopted in sampling for ADV’s following and overtaking, we only capture nearby agents and assume they will interact with the ego. In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas, filtering out the candidates with higher results as the target prior. Those landmarks indicate areas where interaction is most likely to happen. The agents’ trajectories will be decoded conditioned on that lane prior:
@@
Here, $f_{agent}$ is the instance encoding as illustrated in Sec. 3.2, $f_{lane}$ is the corresponding lane encoding obtained in the same way. For the training phase, the ground truth $\tilde{lane}$ corresponding to the agent’s future trajectory will be labeled and incorporated as a condition prior:
@@ 3.5 Trajectory Evaluation
The previous section described a method for multi-stage prediction that outputs grid maps and interactive trajectories. We now show that such grid maps combined with perception occupancy can be integrated directly into an end-to-end motion planner that makes use of space-time cost volumes for scoring candidate trajectories. We follow[48] in max-margin planning, but modify their derivation to take into account scene occupancy and predicted grid map. Our evaluator decodes a space-time cost volume $C_{xy}^{t}=dec(f_{ras})$ in grid form over future timestamps given BEV raster. The cost of a trajectory is defined as the sum of costs at its space-time way-points. The best candidate trajectory is the one with the lowest cost according to the learned cost:
@@
where $S$ represents the set of viable future trajectories, $O_{cc}$ represents perception occupancy, $G_{rid}$ is the predicted grid map, $\alpha$ and $\beta$ are the weights respectively. Since selecting the minimum-cost trajectory within a discrete set is not differentiable, we use the max-margin loss to penalize trajectories that have small cost and are different from the human driving trajectory or are unsafe. Let $\hat{s}$ and $\tilde{s}$ be the input and human trajectory respectively for a given example. We utilize the max-margin loss to encourage the human driving trajectory to have a smaller cost than other trajectories.
@@
The $Diff$ value here serves as compensation to quantitatively measure the difference between the human demonstrated and our candidates. It has to tell how goodness of the human driving or how badness of our samplers, quantified by planning targets, avoiding static collisions, and managing dynamic interactions, etc. So that the evaluator can assign comparable costs to the trajectories in our candidate set. In particularly:
$Diff(\tilde{s},\hat{s})=L2(\tilde{s},\hat{s})+O(\hat{s})+G(\hat{s})$ (6)
@@ 3.6 Trajectory Generation
Given the learned cost volume, the final trajectory can then be computed by minimizing Eq. 4. Note, however, that this optimization with physical constraints is NP-hard. On one hand, we evaluate and extend previous works on trajectory sampling[48, 3, 38, 21] with a more sophisticated design by incorporating prior knowledge and structured outputs from perception and prediction. Furthermore, we leverage a GAN-based pipeline during the generator’s training to solve the optimization problem. In this section, we describe our sampling and generation algorithm.
The sampling and generating algorithms can be mainly categorized into reference line-based and reference line-free. Methods incorporating reference information can generate long-term consistent outputs in the loops. Since this information rarely changes over time, bridging consecutive frames promotes cross-frame stability. However, scenarios still need to be handled where reference-related information can hardly be obtained. To enhance the scalability of our system, we follow [48, 3] to build curve and retrieval-based sampler in reference-line free and illustrate them in the appendix. The proposed lattice sampler, imitation planner, and GAN-based planner will be explained in this section. They incorperate reference info constructed by visual landmarks or commuting trajectories.
@@
The lane centers and surrounding agents’ movements are strong priors to construct the potential trajectory to be executed by the ADV. When landmarks are available, the lane geometry can be exploited to guide the trajectory sampling process. The predicted trajectory, on the other hand, can be incorporated to generate interactive behavior like following or overtaking in longitude. In our paradigm, they are both available as depicted by Sec. 3.4 Sec. 3.1. In particular, we follow the trajectory parameterization and sampling philosophy proposed in [45, 39]. Our sampled trajectories consist of lane keep, lane change, nudge, side pass in lateral, and follow, overtake, yield, stop in longitude. We illustrate the detail in the appendix.
@@
For the imitation planner, previous approaches choose to mimic a human driver with one single trajectory[22, 24]. But such designs suffer from model collapse in both longitude and lateral. Owing to our generator-evaluator paradigm, we can generate a bunch of diverse candidates and pass them to the downstream evaluation task. We follow multiple-trajectory[8] design in trajectory prediction to estimate the conditional distribution $P(S|f_{plan})$ in motion planning task. Where $S=[\hat{s_{1}},...\hat{s_{m}}]$ is the trajectories with $m$ different maneuvers such as lane following, lane changing, agent following, agent overtaking. $f_{plan}$ stands for feature prepared for planning task. ||||| | stands for concatenation.
$f_{plan}=[f_{scene}||f_{route}||f_{ADV}]$ (7)
@@
Our imitation generator decodes a multitude of $M$ possible trajectories for ego move in local coordinates.
@@
During the training phase, we then identify mode $m^{\ast}$ that is closest to the ground-truth trajectory according to distance function,
@@
We select $L2$ dist for the distance function.
GAN Based Planner
The above imitation generator is trained separately from our evaluator by different headers. With the attempt to generate trajectory with minimal cost, we propose a GAN base planner with modification to the training paradigm that bridges the isolated optimization targets, so that Eq. 4 can be solved by learning. For standard adversarial loss[16], also known as the min-max loss, the generator tries to minimize the following function while the discriminator tries to maximize it.
@@
$log(D(x))$ refers to the probability that the discriminator is rightly classifying the real case, maximizing $log(1-D(G(z)))$ would help it to label the fake sample that comes from the generator correctly. In our paradigm, we replace the discriminator with an evaluator. But unlike the standard discriminator in GAN, which gives qualitative labels that indicate true or false, our evaluator outputs a score to measure the discrepancy between candidates quantitatively. So the optimization target for evaluator Eq. 5 still holds, but the learning target of the imitation planner becomes:
$\ L_{generator}=E(G(m))+Diff(\tilde{s},G(m))$ (12)
@@
$E(s)=C(s)=\sum_{(x_{t},y_{t})\in{s}}C(x_{t},y_{t})$ (14)
$\ C_{xy}^{t}=dec(f_{ras})$ (15)
$\ G(m)=\hat{S}=[\hat{s_{1}},...\hat{s_{m}}]$ (16)
Here, m is a hidden variable representing the manual in planning.
However, the trajectory evaluation Eq. 14 involve index trajectory points in the learned cost volume $C_{xy}^{t}=dec(f_{ras})$, which is non-differentiable, thus cannot be back propagated to the generator. To address this issue, we use a soft function to get a meaningful gradient. The grid_sample operator[1] is adopted here to obtain the gradients via interpolation between values in our cost volume.
@@ 3.7 Safety layer
For on-road autonomous driving motion planning, a strategy for safety coverage is necessary. The perception occupancy and trajectory prediction results are used to validate the trajectory candidates with respect to safety. The trajectory prepared for the safety layer is a little bit different to Eq. 2 in that no prior is assumed in decoding. We sort the trajectories according to the evaluated costs and iterate them to check the basic rule in modularized form(like trajectory and bounding box) until we cache enough candidates by ascending order. To ensure cross-frame consistency, we reserve $n$ validated trajectories with the lowest cost and compare them with planning output in the last frame. The above objective promotes the system’s robustness with respect to the corner cases that our model fails to handle. We believe such a layer of safety guarantee is indispensable for manufacturing our system at scale.
@@ 4 Experiments
We develop various baseline variants of our models under both open-loop and closed-loop configurations for comprehensive ablative analyses, and we quantitatively benchmark these against cutting-edge alternative methodologies. We gather in-house datasets for both training and open-loop evaluations, encompassing an extensive array of driving scenarios within densely populated urban settings. To substantiate the feasibility of the data-driven planners, we deactivate the functionalities in the safety layer as well as the optimizer and safety backup subsystems in comparative methodologies such as optimal control in UniAD[22], etc.
@@ 4.1 Prediction
Method Mean L2 / hit-rate $<$ 1m Avg.
1s 2s 3s
VectorNet[13] 0.75/0.80 1.72/0.40 2.85/0.22 1.34
GAD-uni 0.76/0.72 1.70/0.38 2.83/0.23 1.35
TNT[50] 0.65/0.78 1.01/0.57 1.72/0.34 0.90
GAD-multi 0.60/0.84 0.95/0.65 1.59/0.39 0.81
Table 1: Prediction Results in uni-modal and multi-modal.
Method trajectory sources L2(m) Collision(%)
curve retrieval lane lattice model 1s 2s 3s 1s 2s 3s
UniAD[22] $\surd_{uni}$ 0.14 0.44 0.92 0.17 0.33 0.78
VAD[24] $\surd_{uni}$ 0.12 0.41 0.89 0.15 0.30 0.69
NMP[48] $\surd$ 0.25 1.19 2.96 0.08 0.16 1.05
MP3[3] $\surd$ 0.42 1.31 2.73 0.09 0.22 0.97
P3[38] $\surd$ 0.24 1.55 3.95 0.14 0.35 1.04
GAD-multi $\surd_{multi}$ 0.16 0.69 1.26 0.11 0.31 0.52
GAD-ref_free $\surd$ $\surd$ 0.25 0.90 2.15 0.09 0.17 0.48
GAD-base $\surd$ $\surd_{multi}$ 0.20 0.84 1.84 0.10 0.31 1.07
GAD-gan $\surd$ $\surd_{gan}$ 0.18 0.80 1.67 0.09 0.28 1.02
Table 2: Open loop planning results of imitation based method, sampling based method and GAD.
Method comfort event in 30Km efficiency MPI
Lat. accel. Jerk $\Delta$Steering Angle Collisions Lost control Discomfort Deviation Minute
UniAD[22] 0.351 3.12 3.68 - - - - 15.20 0.31
ST-P3[21] 0.857 4.54 8.53 4 3 14 3 8.62 1.36
GAD-multi 0.489 1.22 2.89 4 1 2 4 5.60 2.72
GAD-lattice 0.728 2.26 4.22 2 0 4 2 5.68 3.75
GAD-base 0.531 1.20 4.21 3 1 3 1 5.56 3.75
GAD-GAN 0.511 1.15 3.52 3 1 2 1 5.51 4.28
DPQP 0.284 0.85 2.11 1 0 1 1 5.42 10.0
Table 3: Road Test Outcomes: Data were systematically collected during a 30 km road test, encompassing the computation of average time per loop. Concurrently, measurements of lateral acceleration, jerk, and steering angle were meticulously recorded. The UniAD consistently deviates and collides with obstacles, thereby failing to accumulate sufficient statistical data.
@@ 4.2 Open-loop Planning
4.3 Closed-loop Experimental

### v3

@@ 3.4 Prediction
We employ a multi-stage design for scene and agent prediction as depicted in Fig. 3, where the first stage, involves predicting the scene-level grid map (also known as the heat map or occupancy) by regressing $G_{rid}=dec(f_{ras})$ with a resolution that matches the rasterization input through decoding of the scene embeding. Temporal resolution $t$ is 0.1 second from 0 to 3 seconds.
The grid prediction decoder, as mentioned in Sec. 3.2, shares a similar structure with the cost volume decoder, yet it is supervised from the rasterization of the agents’ future ground truth box. This grid is trained with pixel-wise focal loss inspired from [16], for brief we replace $G_{rid}$ with $g$:
$\begin{array}[]{r}L_{pred\_grid}=-\sum_{ijt}[\tilde{g_{ijt}}(1-\hat{g_{ijt}})^{2}log(\hat{g_{ijt}})+\\ (1-\tilde{g_{ijt}})\hat{g_{ijt}}^{2}log(1-\hat{g_{ijt}})]\end{array}$ (1)
Here, $\tilde{g_{ijt}}$ represents the class label for the grid, where $\tilde{g_{ijt}}=1$ indicates that it is occupied in timestamp t. In the second stage, the model forecasts agent-level trajectories with the scene prior from the predicted grid by following an interactive design. Since the output trajectories are mainly adopted in sampling for ADV’s following and overtaking, it captures only nearby agents, assuming interactions with the ego. In particular, we gather the corresponding probability of nearby landmarks by accumulating grid probability belonging to those areas $P_{lane}=\sum_{t}\sum_{i,j\in(lane)}g_{ij}^{t}$, filtering out higher probability candidates as the target prior. These landmarks indicate likely interaction areas, and agents’ trajectories are decoded based on this lane prior:
@@
Here, $f_{agent}$ is the instance encoding as illustrated in Sec. 3.2, $f_{lane}$ is the corresponding lane encoding obtained in the vectorized stream. For the training phase, the ground truth $\tilde{lane}$ corresponding to the agent’s future trajectory will be labeled and incorporated as a condition prior:
@@ 3.5 Trajectory Evaluation
The previous section introduced a multi-stage prediction method that produces grid maps and interactive trajectories. We now demonstrate the grid can be seamlessly integrated with perception occupancy into an end-to-end motion planner utilizing space-time cost volumes to evaluate candidate trajectories. Following the framework of [51] for max-margin planning, we refine it to include scene occupancy and anticipated grid maps, and our evaluator decodes a space-time cost volume $C_{xy}^{t}=dec(f_{ras})$ in grid form for future 30 consecutive frame from BEV raster. The trajectory cost is calculated as the sum of costs at its space-time waypoints, with the best trajectory being the one that minimizes this cost:
@@
where $S$ represents the set of planed trajectories, $O_{cc}$ represents perception occupancy, $G_{rid}$ is the predicted grid map, $\alpha$ and $\beta$ are hyperparameters. Since selecting the minimum-cost trajectory within a discrete set is not differentiable, we use the max-margin loss to penalize trajectories that either deviate from the human driving trajectory or are unsafe, and to encourage the human driving trajectory $\tilde{s}$ to have a smaller cost than other trajectories $\hat{s}$ for a given example:
@@
The $Diff$ metric functions as a compensatory quantifier, meticulously gauging the disparity between human demonstrations and algorithmic candidates. It is designed to assess the efficacy of human driving and the deficiencies of candidates, as delineated by objectives such as planning target, evasion of static obstacles, and regulation of dynamic interactions. This enables the evaluator to attribute commensurate costs to the given trajectories. In particularly:
$Diff(\tilde{s},\hat{s})=L2(\tilde{s},\hat{s})+O(\hat{s})+G_{rid}(\hat{s})$ (6)
@@ 3.6 Trajectory Generation
Given the learned cost volume, the final trajectory can then be computed by minimizing Eq. 4. Note, however, that this optimization with physical constraints is NP-hard. On one hand, we evaluate and extend previous works on trajectory sampling[51, 3, 41, 22] by incorporating prior knowledge and structured outputs from perception and prediction. Furthermore, we leverage a GAN-based pipeline during the imitation generator’s training to solve the optimization problem. In this section, we describe our sampling and generation algorithm.
The sampling and generating algorithms are categorized into reference line-based and reference line-free types, where methods that incorporate reference information provide long-term consistent outputs and enhance cross-frame stability as this information remains largely unchanged over time. However, challenges arise in scenarios where reference-related information is scarce. To improve our system’s scalability, we follow [51, 3] to develop curve and retrieval-based samplers in a reference-line free context, detailed in Sec. A.8 and Sec. A.9. Additionally, this section will discuss the proposed lattice sampler, imitation planner, and GAN-based planner, which utilize reference information constructed from visual landmarks or commuting trajectories.
@@
The lane centers and surrounding agents’ movements are strong priors to construct the potential trajectory to be executed by the ADV. When landmarks are available, the lane geometry can be exploited to guide the trajectory sampling process. The predicted trajectory, on the other hand, can be incorporated to generate interactive behavior like following or overtaking in the longitudinal direction. In our paradigm, they are both available as depicted by Sec. 3.4 Sec. 3.1. In particular, we follow the trajectory parameterization and sampling philosophy proposed in [48, 42]. The sampled trajectories include maneuvers such as lane keeping, lane changing, nudging, side passing in the lateral direction, as well as following, accelerating, yielding, and stopping in the longitudinal direction. More details can be found in Sec. A.10.
@@
For the imitation planner, previous approaches choose to mimic a human driver with one single trajectory[23, 25]. But such designs suffer from model collapse in both longitudinal and lateral directions. Owing to our generator-evaluator paradigm, we can generate a bunch of diverse candidates and pass them to the downstream evaluation task. We follow multiple-trajectory[8] design to estimate the conditional distribution $P(S|f_{plan})$ and multi-modality in motion planning task:
@@
Where $S=[\hat{s_{1}},...\hat{s_{m}}]$ is the trajectories with $m$ different maneuvers such as lane following, lane changing, agent following, agent overtaking. $f_{plan}$ is constructed by feature prepared for planning task in Eq. 8, where (||||| |) stands for concatenation.
$f_{plan}=[f_{scene}||f_{route}||f_{ADV}||f_{target\_speed}]$ (8)
The imitation generator decodes $M$ possible trajectories for ego move in local coordinates. During training, we then identify modal closest to the ground-truth trajectory according to $L2$ distance.
@@
GAN-Based Planner
The aforementioned imitation generator is trained independently from our evaluator under distinct supervisory decoder. In pursuit of generating trajectories utilizing the acquired cost volume and addressing Eq. 4 through a learning-based approach, we propose a GAN-based planner with a conditional generative adversarial training framework. This framework is designed to bridges the isolated optimization targets between generator and evaluator, thus facilitating the learning of a stochastic generative model beyond trajectory imitation, adept at encapsulating the multi-modal diversity inherent in planning trajectories. Regarding the standard adversarial loss[17], commonly referred to as the min-max loss, the generator endeavors to minimize the subsequent function, while the discriminator seeks to maximize it.
@@
$log(D(x))$ refers to the probability that the discriminator is rightly classifying the real case, maximizing $log(1-D(G(z)))$ would help it to label the fake sample that comes from the generator correctly. In our proposed framework, the conventional discriminator is supplanted by the evaluator as delineated in Sec. 3.5. Contrary to the traditional discriminator in GANs, which categorizes instances with qualitative labels denoting veracity, our evaluator quantifies the divergence among candidates through the learned cost volume. Thereby, the optimization target for evaluator Eq. 5 still holds, but the learning target of the generator changed. The complete formula derivation is provied in Sec. A.11.
$\ L_{generator}=E(G(m))+L_{imit\_plan}(\tilde{s},G(m))$ (12)
@@
$E(s)=\sum C(s)+\alpha O_{cc}(s)+\beta G_{rid}(s)$ (14)
$[\hat{s_{1}},...\hat{s_{m}}]=\hat{S}=G(m);m\sim\mathcal{N}(\mu,\,\sigma^{2})\,.$ (15)
The architecture is exactly the same as presented in previous subsections, except that in the generative model, the hidden variable $m$, which indidates driving manual in deterministic model is sampled from a Gaussian distribution. The final encoding for ADV is concatenated with $m$ to decode into its trajectory. However, the trajectory evaluation Eq. 14 involve index trajectory points in the learned cost volume $C_{xy}^{t}=dec(f_{ras})$, which is non-differentiable, thus cannot be back propagated to the generator. To address this issue, we use a soft function to get a meaningful gradient. The grid_sample operator[1] is adopted here to obtain the gradients via interpolation between values in our cost volume.
@@ 3.7 Safety layer
For on-road autonomous driving, a strategy for safety coverage is necessary. The perception occupancy and trajectory prediction results are used to validate the trajectory candidates with respect to safety. The trajectory prepared for the safety layer is a little bit different to Eq. 2 in that no prior is assumed in decoding. We sort the trajectories according to the evaluated costs and iterate them to check the basic rule in modularized form(like trajectory and bounding box) until we cache enough candidates by ascending order. To ensure cross-frame consistency, we reserve $n$ validated trajectories with the lowest cost and compare them with planning output in the last frame. The above objective promotes the system’s robustness with respect to the corner cases that our model fails to handle. We believe such a layer of safety guarantee is indispensable for manufacturing our system at scale.
@@ 4 Experiments
We develop various baseline variants of our models under both open-loop and closed-loop experiments for comprehensive ablative analyses, and we quantitatively benchmark these against cutting-edge alternative methodologies. We gather in-house datasets for both training and open-loop evaluations, encompassing an extensive array of driving scenarios within densely populated urban settings. To substantiate the feasibility of the data-driven planners, we deactivate the functionalities in the safety layer as well as the optimizer and safety backup subsystems in comparative methodologies such as optimal control in UniAD[23], etc.
@@ 4.1 Prediction
Method Mean L2 / hit-rate $<$ 1m Avg. 1s 2s 3s VectorNet[14] 0.75/0.80 1.72/0.40 2.85/0.22 1.34 GAD-uni 0.76/0.72 1.70/0.38 2.83/0.23 1.35 TNT[53] 0.65/0.78 1.01/0.57 1.72/0.34 0.90 GAD-multi 0.60/0.84 0.95/0.65 1.59/0.39 0.81
Table 1: Prediction Results in uni-modal and multi-modal.
@@ 4.2 Open-loop Planning
Method hist tar v L2 passing rate
ras vec m/s Ave straight turn
GAD-multi_1 $\surd$ $\surd$ x 0.56 0.00 0.00
GAD-multi_2 $\surd$ $\surd$ $\surd$ 0.63 0.80 0.75
GAD-multi_3 $\surd$ x $\surd$ 0.70 0.85 0.70
GAD-multi_4 x x $\surd$ 0.34 0.65 0.00
Table 2: Ablation study on ego feature for imitation planner.
Method trajectory sources L2(m) Collision(%) curve retrieval lane lattice model 1s 2s 3s 1s 2s 3s UniAD[23] $\surd_{uni}$ 0.14 0.44 0.92 0.17 0.33 0.78 VAD[25] $\surd_{uni}$ 0.12 0.41 0.89 0.15 0.30 0.69 NMP[51] $\surd$ 0.25 1.19 2.96 0.08 0.16 1.05 MP3[3] $\surd$ 0.42 1.31 2.73 0.09 0.22 0.97 P3[41] $\surd$ 0.24 1.55 3.95 0.14 0.35 1.04 GAD-multi $\surd_{multi}$ 0.16 0.69 1.26 0.11 0.31 0.52 GAD-ref_free $\surd$ $\surd$ 0.25 0.90 2.15 0.09 0.17 0.48 GAD-base $\surd$ $\surd_{multi}$ 0.20 0.84 1.84 0.10 0.31 1.07 GAD-gan $\surd$ $\surd_{gan}$ 0.18 0.80 1.67 0.09 0.28 1.02
Table 3: Open loop planning results of imitation-based method, sampling-based method and GAD.
4.3 Ablation of Ego Feature
A potentially misleading aspect of current data-driven methodologies in motion planning may be the excessive reliance in utilization of ego history. Historical way points are pivotal in predictive tasks for inferring the future trajectories of agents. Conversely, planning tasks diverge in that the trajectory is synthesized to achieve specific longitudinal and lateral objectives such as velocity and lateral position, thereby necessitating minimal dependence on historical motion. We conduct experiments to validate the necessity of ego history and target speed encoding in Tab. 2. The historical information is derived from two distinct sources within our framework: one pertaining to rasterized ego motion and the other to vectorized historical positions. Since lateral targets can be deduced from route and landmarks, we only provides feature encoding in target speed of 3s-th to guide trajectory generation(it can be conceptualized as a data-driven sampler analogous to the cruise scenario strategy within lattice framework but only concerns the speed in 3s-th). Such setting provides a tunable parameter in deployment for the users to adjust the desired cruise speed range. As show in the Tab. 2, we evaluate the open-loop metric and the closed-loop scene passing rate of lane keeping and turning for imitation planner. The results indicate that history feature dose help to improve the open-loop metric, but target speed feature is the key for success closed-loop. Method without history encoding failed in turning scene.
4.4 Closed-loop Experiments